OpenStudy (zenmo):

Determine whether the series is absolutely convergent, conditionally convergent, or divergent.

zepdrix (zepdrix):

So this is a limit problem? Not differentiation. Ok ok

zepdrix (zepdrix):

L'Hopital's rule will involve differentiation :) Yes, good call.

zepdrix (zepdrix):

\[\large\rm y=\lim_{x\to\infty}\left(1+\frac1x\right)^x\]If we try to evaluate this using direct substitution, we get the indeterminate form \(\large\rm 1^{\infty}\). So ummm, ok ya your steps look good so far.

zepdrix (zepdrix):

Sure, if you'd like :)\[\large\rm \ln y = \lim_{n\to\infty}\frac{\ln\left(1+\frac1n\right) }{\frac1n}\]

zepdrix (zepdrix):

\[\large\rm \ln y = \lim_{n\to\infty}\frac{\ln\left(1+t\right) }{t}\]So what happens with our limit? Where is t going as n grows?

OpenStudy (zenmo):

infity over infity, so indeterminate form and use L'Hospital again?

zepdrix (zepdrix):

No, I mean the thing written UNDER the limit notation, see it? :) n is approaching infinity, what's happening with t, not with the entire expression.

OpenStudy (sprattfield):

\[\sqrt{2}\]

OpenStudy (mathmale):

Zenmo: I'm curious: why doesn't your math problem appear in your post? And what about the instructions for this problem (find what...?... using log diff'n.)?

OpenStudy (mathmale):

My comment (above) pertained to the previous problem that you posted, but my questions are still valid.

OpenStudy (zenmo):

\[\sum_{n=1}^{\infty}(1+\frac{ 1 }{ n })^{n^2}\] Original Problem. Yes, give me time to type it all out and you will see my point (of question).

OpenStudy (mathmale):

Your "Determine whether the series is absolutely convergent, conditionally convergent, or divergent." does not apply to the expression you typed in:\[\lim_{n \rightarrow \infty}(1+\frac{ 1 }{ n })^n\]

OpenStudy (mathmale):

I'd suggest you start over. Make sure that your instructions and your math statement match.

OpenStudy (sprattfield):

hi mathmale >:)

OpenStudy (zenmo):

\[\left| a _{n} \right|=a _{n}=(1+\frac{ 1 }{ n })^{n^2}\] Using "RATIO TEST," \[\lim_{n \rightarrow \infty}\sqrt[n]{(1+\frac{ 1 }{ n })^{n^2}}=\lim_{n \rightarrow \infty}(1+\frac{ 1 }{ n })^{\frac{ n^2 }{ n }}=\lim_{n \rightarrow \infty}(1+\frac{ 1 }{ n })^n\]

OpenStudy (zenmo):

Now, I need to find the limit (which I need to apply differentiation).

zepdrix (zepdrix):

Did you understand my last comment? We didn't finish making our substitution,\[\large\rm \ln y = \lim_{\color{red}{n\to\infty}}\frac{\ln\left(1+t\right) }{t}\]

OpenStudy (zenmo):

Yes, but mathdale wanted to see my question.

zepdrix (zepdrix):

ya np :)

OpenStudy (zenmo):

does ln(infinity) approaches infinity ?

zepdrix (zepdrix):

Yes, but maybe you're still confused. We're not allowed to plug anything in for t until you fix this notation under the limit.

OpenStudy (zenmo):

Oh the format is wrong.

OpenStudy (zenmo):

\[\ln y = \lim_{t \rightarrow \infty}\frac{ \ln(1+t) }{ t }\]

zepdrix (zepdrix):

\(\large\rm n\to\infty\) so then if \(\large\rm t=\frac1n\), we have \(\large\rm t\to?\)

zepdrix (zepdrix):

No, but good guess! :)

zepdrix (zepdrix):

Notice that t is the flip of n, right?

OpenStudy (zenmo):

t goes to infinity?

zepdrix (zepdrix):

(This is one of the reasons I don't like the substitution, hehe)

zepdrix (zepdrix):

No. As n goes to infinity, the reverse of that will be, t goes toward zero.

zepdrix (zepdrix):

\[\large\rm n\to\infty:\qquad \frac{1}{n}\to0\]

OpenStudy (zenmo):

Yea, with substitution it would be infinity over infinity. Without sub, it will be 0 over 0. Yes?

zepdrix (zepdrix):

Yes, good call :D You're missing some of the important details, but it will still give us the indeterminate form that we're looking for.\[\large\rm \ln y = \lim_{t\to0}\frac{\ln\left(1+t\right) }{t}\]

zepdrix (zepdrix):

It's giving us the indeterminate form that we need in order to be ALLOWED to apply L'Hopital's Rule: 0/0.

OpenStudy (zenmo):

ahh ... yes

OpenStudy (zenmo):

Yea, forgot that the limit as "some variable" need to --> 0 instead of infinity when using LH.

zepdrix (zepdrix):

No fancy quotient rule or anything like that, which is nice. We differentiate the top and bottom `separately`.

zepdrix (zepdrix):

Remember derivative of your log function? :d

OpenStudy (zenmo):

\[\ln(1+t) = \frac{ 1 }{ t }\]

OpenStudy (zenmo):

Denominator becomes 1, so 1 / t after differentiating?

zepdrix (zepdrix):

\[\rm \frac{d}{dt}\ln(1+t)=\frac{1}{1+t}\]

OpenStudy (zenmo):

Oh yea, gotcha

zepdrix (zepdrix):

\[\large\rm \ln y = \lim_{t\to0}\frac{\ln\left(1+t\right) }{t}\]Hmm ok that cleans things up pretty nicely,\[\large\rm \ln y = \lim_{t\to0}\frac{1}{1+t}\]

zepdrix (zepdrix):

So what next? Maybe direct substitution?

OpenStudy (zenmo):

Sorry, post bugged out earlier. I got the problem now. Thanks zepdrix :)